im trying to solve this problem showing step by step; nP6=5(nP5)it is alegebra 2
Answer by jim_thompson5910(35256)

n P 6 = n(n-1)(n-2)(n-3)(n-4)(n-5)


n P 5 = n(n-1)(n-2)(n-3)(n-4)

-------------------------------

n P 6 = 5*( n P 5 )

n(n-1)(n-2)(n-3)(n-4)(n-5) = 5 * [ n(n-1)(n-2)(n-3)(n-4) ]

n-5 = 5 ... divide both sides by n(n-1)(n-2)(n-3)(n-4),

this is last step only done if n doesn't equal 0, 1, 2, 3, or 4.

n = 5+5

n = 10

So the answer is n = 10